Westwood Grocery Store needs a new freezer for popsicles and ice cream bars. The owner finds a freezer that is 4 feet long, 2 feet wide, and 3 feet tall. What is the volume of the freezer?

1
Expert's answer
2022-03-14T10:57:12-0400

Volume "=Length\times width\times height"

Volume = "4ft\times 2feet\times 3feet"

Volume = "24 ft^3"


"V=24ft^3"